OpenStudy (madhu.mukherjee.946):

a loan earns interest at a certain rate (compound)per cent per annum.four years ago the amount was rs.81 ;now it is rs.144.what will the amount be two years hence?

OpenStudy (anonymous):

Use the formula and plug in the values.

OpenStudy (anonymous):

\[A=P(1+\frac{ r }{ 100 })^n\]

OpenStudy (madhu.mukherjee.946):

dude i know the formula but this sum is different

OpenStudy (madhu.mukherjee.946):

basically we neither have p value nor t nor r only we have is two compound interest amount

OpenStudy (anonymous):

you have P, A and n

OpenStudy (madhu.mukherjee.946):

where do you see p and 81 and 144 are compound interest amount and not the total amount

OpenStudy (anonymous):

it is the interest+year 0 p

OpenStudy (madhu.mukherjee.946):

u solve the problem if you can and pls give me the steps coz i can't

OpenStudy (anonymous):

first rearrange the formula and put r to one side

OpenStudy (anonymous):

\[144=81(1+\frac{ r }{ 100 })^4\] \[\frac{ 144 }{ 81 }=(1+\frac{ r }{ 100 })^4\]

OpenStudy (anonymous):

\[\frac{ 16 }{ 9 } = (1+\frac{ r }{ 100 })^4\]

OpenStudy (anonymous):

\[\frac{ 4 }{ 3 } = (1 + \frac{ r }{ 100 })^2\]

OpenStudy (anonymous):

\[\frac{ 4 }{ 3 } = 1+ \frac{ r }{ 50 } + \frac{ r^2 }{ 10000 }\]\[-\frac{ 1 }{ 3 } = \frac{ r }{ 50 } + \frac{ r }{ 10000 }\]

OpenStudy (madhu.mukherjee.946):

i do not think this is ok...coz the answer given id 191.40

OpenStudy (anonymous):

some typos there

OpenStudy (madhu.mukherjee.946):

ok thanks anyways

OpenStudy (anonymous):

mines 15.47

OpenStudy (madhu.mukherjee.946):

no its 191.40

OpenStudy (madhu.mukherjee.946):

thanks for you try

OpenStudy (anonymous):

the interest rate is 15.47

OpenStudy (anonymous):

so the 2 years stuff its up to you to try

OpenStudy (madhu.mukherjee.946):

i'm so sorry saseal pls don't mind your calculation was absolutely correct .....i understood the sum a little later thanks dude thank you so much